x-2y=5 and 4x+y=65
Answer by ewatrrr(24785)


Hi,
Solving by elimination of the y variable
x-2y=5 and
4x+y=65
x -2y = 5
8x+2y= 130 mulitplying 2nd equation by 2 to eliminate y
9x = 135
x = 15
solving for y
x - 2y = 5
15 - 2y = 5
10 = 2y
5 = y
CHECKING our Answer
x-2y=5
15 - 2*5 = 5